OpenStudy (anonymous):

Helium is pumped into a spherical balloon at a constant rate of 4 cubic feet per second. How fast is the radius increasing after 3 minutes.At what time is the radius increasing at a rate of 100 feet per second?

jimthompson5910 (jim_thompson5910):

V = (4/3)*pi*r^3 dV/dt = (4/3)*3*pi*r^2*dr/dt

jimthompson5910 (jim_thompson5910):

dV/dt = (4/3)*3*pi*r^2*dr/dt dV/dt = 4*pi*r^2*dr/dt

OpenStudy (anonymous):

ok i got that

OpenStudy (anonymous):

PLEASE HELP ME IM REALLY STUCK!!!!!!!!!

OpenStudy (anonymous):

The volume of a sphere is: V = (4/3)πr^3. By implicit differentiation: dV/dt = 4πr^2(dr/dt). We know dV/dt and wish to find dr/dt. In order to find dr/dt, we also need to know r. After 3 minutes. Notice that 3 minutes = 180 seconds. Then, after 3 minutes, the volume of the balloon is 180(3) = 540 cubic feet. At this point: 540 = (4/3)πr^3, by substituting V = 540 ==> r = (405/π)^(1/3). Therefore: dV/dt = 4πr^2(dr/dt) ==> dr/dt = (dV/dt)/(4πr^2) = 3/[4π(405/π)^(2/3)] ≈ 0.0094 ft/s.

OpenStudy (anonymous):

Hope this helps.

OpenStudy (anonymous):

but i have an answer and its telling me it is 0.0103 ft./sc and i dont know how to get it

OpenStudy (anonymous):

Just follow the formula I gave you and you should find it.

OpenStudy (anonymous):

is there any other info... like r(t) = ???

OpenStudy (anonymous):

double check to see if they give a formula for the radius, r(t) = ???

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you have to find r

OpenStudy (anonymous):

helium is pumped into a spherical balloon ar a constant rate of 4 cubic feet per second. How fast is the radius increasing after 3 minuites?

OpenStudy (anonymous):

at what time(if any) is the radius increasing at a rate of 100 ft per second?

OpenStudy (radar):

I believe that the 180(3) in the @TurtleNadz post above (the long post) should be 180(4) for 720 for volume, and then calculate r

OpenStudy (radar):